Please help me solve this equation:
Solve using the elimination method. Also determine whether each system is consistent or inconsistent and whether the equations are dependent or independent.
3x-12y=6
2x-8y=4


Thank you for your time I really appreciated

Answer by TimothyLamb(4379)

3x-12y=6
2x-8y=4
---
put the system of linear equations into standard form
---
x + y = 9
5x + 10y = 65
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
This linear system is singular. It has no solution
---
dependent and inconsistent
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---